OpenStudy (selena2345):

The ratio of men to women at a class is 6 to 5. How many women students are there if there are 3600 men ?

OpenStudy (texaschic101):

proportion time ... men / women : 6/5 = 3600/x --- 6 men to 5 women = 3600 men to x women cross multiply (6)(x) = (5)(3600) 6x = 18000 x = 18000/6 x = 3000 check.. 6/5 = 3600/3000 cross multiply (6)(3000) = (5)(3600) 18000 = 18000 (correct) there will be 3000 men when there is 3600 women